OpenStudy (anonymous):

Help? Please and thank you! Two stores sell the same computer for the same original price. Store A advertises that the computer is on sale for 25% off the original price. Store B advertises that it is reducing the computer’s price by $180. When Brittany compares the sale prices of the computer in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the computer’s original price.

OpenStudy (anonymous):

I'll give a medal.......

OpenStudy (anonymous):

and a fan?????????

OpenStudy (anonymous):

Perhaps....

OpenStudy (anonymous):

k ill look at it

OpenStudy (anonymous):

So kind.

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

cant figure out which equation to use

OpenStudy (anonymous):

Do you want the options?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

A. 0.25p = p – 180 B. 0.75p = p + 180 C. 0.25(p – 180) = p D. 0.75p = p − 180

OpenStudy (anonymous):

Oops.

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

aight

OpenStudy (anonymous):

Eh, My computer is acting stupid. Can you see the options?

OpenStudy (anonymous):

yea r those the right ones?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Any guesses?

OpenStudy (anonymous):

i think its A. not sure though

OpenStudy (anonymous):

Alright thanks man.

OpenStudy (anonymous):

cya

OpenStudy (anonymous):

@cody7744 Have an opinion?

OpenStudy (anonymous):

it isA i know it for a fact

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

your welcome